OpenStudy (anonymous):

If 7x is 24 more than X, then X^2 is how much more than sqrt x.

OpenStudy (anonymous):

A. 49 - sqrt7 B. 24 C. 14 D. 2sqrt6 E. 6

OpenStudy (anonymous):

Write the first half as an equation, solve for x. Then plug x into the 2 questions in the 2nd half and take the difference

OpenStudy (anonymous):

So, 7x = 24 +x (subtract x from both sides) 6x = 24 (divide both sides by 6) x = 4 x^2 = 4^2 = 16 sqrt x = sqrt 4 = 2 16 -2 = 14

OpenStudy (phoenixfire):

Well, he wrote x and X, so could it be that they're two different variables?

OpenStudy (anonymous):

no same variable...in the question is says the 7x is 24 more then x. That's what my equation says..... 7x = 24 + x ...plug in x to confirm 7(4) = 24 + 4 28 = 28

OpenStudy (anonymous):

No it's only one it's just my iPad corrected the other x to s capital

OpenStudy (anonymous):

oh sorry @PhoenixFire, misunderstood your question